Banyan Co.’s common stock currently sells for $54.25 per share. The growth rate is a constant 6%, and the company has an expected dividend yield of 6%. The expected long-run dividend payout ratio is 20%, and the expected return on equity (ROE) is 7.5%. New stock can be sold to the public at the current price, but a flotation cost of 5% would be incurred. What would be the cost of new equity? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Expected Dividend Yield = D1/Stock Price

So,

D1 = 0.06(54.25) = $3.255

Using DDM Model,

Stock Price(1 - f) = D1/(r - g)

54.25(1 - 0.05) = 3.255/(r - 0.06)

r = 0.1232

Cost of New Equity = 12.32%